OpenStudy (hang254):

What is the output of the given input f(x) = -x^2-x, find f(-4)

OpenStudy (anonymous):

just plug in -4 into everywhere where there is an x

OpenStudy (anonymous):

so solve -4^2 - 4 = ?

OpenStudy (hang254):

F(x) = 12?

OpenStudy (anonymous):

careful here

OpenStudy (anonymous):

dude you are top dog man

OpenStudy (anonymous):

\(-x^2\) means to square \(x\) and then put a minus sign in front

OpenStudy (hang254):

So -4

OpenStudy (anonymous):

oh yeah so its +4

OpenStudy (anonymous):

\[f(x) = -x^2-x\] \[f(-4)=-(-4)^2-(-4)\]

OpenStudy (hang254):

I mean -2

OpenStudy (anonymous):

good call satellite

OpenStudy (anonymous):

be very careful, you have not posted the correct answer yet

OpenStudy (anonymous):

\[f(-4)=-(-4)^2-(-4)=-(-4)(-4)+4=-16+4=...\]\]

OpenStudy (hang254):

Hmmmm, f(x)= 20

OpenStudy (hang254):

@satellite73

OpenStudy (anonymous):

\(-16+4=4-16\neq 20\)

OpenStudy (hang254):

-12= -12 = 20 ?

OpenStudy (hang254):

@Chlorophyll

OpenStudy (anonymous):

f ( -4 ) = - ( -4)² - ( -4) = - 16 + 4 = ...?

OpenStudy (hang254):

-12

OpenStudy (hang254):

Thanks so much :)

OpenStudy (anonymous):

Just plug in exactly the way I show you, so you won't miss the sign :)

OpenStudy (hang254):

Thanks, I will next time